Jeremiah can kayak downstream a distance of 21 kilometers in 3 hours. Going upstream, he can kayak a distance of only 24 kilometers in 4 hours. How fast could Jeremiah kayak in still water (if there were no current)?
Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

Jeremiah can kayak downstream a distance of 21 kilometers in 3 hours. Going upstream, he can kayak a distance of only 24 kilometers in 4 hours. How fast could Jeremiah kayak in still water (if there were no current)?
-----------------------------------
Downstream rate: dist/time = 21/3 = 8 km/hr
--------
Upstream rate: dist/time = 24/4 = 6 km/hr
-----
Equations
downstream Eq: b + c = 8 km/hr
upstream Eq::: b - c = 6 km/hr
------------------------------------
Add and solve for "b":
2b = 14 km/hr
boat speed = 7 km/hr
-----
Solve for "c":
b + c = 8
7 + c = 8
current = 1 km/hr
======================
Cheers,
Stan H.
============

Answer by Alan3354(69443)

Jeremiah can kayak downstream a distance of 21 kilometers in 3 hours. Going upstream, he can kayak a distance of only 24 kilometers in 4 hours. How fast could Jeremiah kayak in still water (if there were no current)?
-------------------------------
d'stream speed = 21/3 = 7 km/hr
u'stream = 24/4 = 6 km/hr
----
His speed is the average = 6.5 km/hr